# Abundance and SYZ conjecture in families of hyperkähler manifolds

### v1

@@ 1.2 The hyperkähler SYZ conjecture
In each of the currently known deformation classes of hyperkähler manifolds there exist manifolds admitting Lagrangian fibrations. Therefore a natural approach to 1.2 (at least for the known deformation classes) is to obtain the result by deformation techniques. This approach was extensively explored by Matsushita, see [Mat3], [Mat4] and [Mat5]. Using a theorem of Voisin [Voi] about deformations of Lagrangian submanifolds of a hyperkähler manifold $M$, Matsushita showed in [Mat4] that if a line bundle $L$ with $q(c_{1}(L))=0$ is semiample, then it remains semiample after a small deformation of the pair $(M,L)$, i.e. semiampleness is open in families.
In [Mat5, Theorem 1.2] Matsushita extends his results and obtains more precise statements about the behaviour of semiampleness under local deformations, assuming additionally that the base of the Lagrangian fibration defined by the semiample line bundle is isomorphic to $\mathbb{C}P^{n}$.
@@ 1.3 Results
In order to prove this result we develop a moduli theory for the pairs $(M,L)$ where $M$ is a hyperkähler manifold and $L$ a semiample line bundle of vanishing BBF square on $M$. Fixing a cohomology class $\ell\in H^{2}(M,{\mathbb{Z}})$, we define a semiample Teichmüller space $\mathcal{T}_{\mathrm{sa}}(M,\ell)$ (see Section 2.4) as a subset of $\mathcal{T}(M,\ell)\subset\mathcal{T}(M)$, where $\mathcal{T}(M,\ell)$ is a divisor in the usual Teichmüller space where the class $\ell$ stays of Hodge type $(1,1)$. We define the period domain for $\mathcal{T}_{\mathrm{sa}}(M,\ell)$ and study the fibres of the corresponding period map. Our main result is a version of the global Torelli theorem for $\mathcal{T}_{\mathrm{sa}}(M,\ell)$, 3.3, claiming in particular that the period map is surjective when the semiample Teichmüller space is non-empty.
Our results rely only on the openness of semiampleness proved by Matsushita in [Mat4] and do not use [Mat5]. In particular, we do not assume anything about the base of our Lagrangian fibrations, which makes our results independent of the well known open problems regarding the smoothness of the base. Our main tool is the construction of special families of Lagrangian fibrations called the degenerate twistor families. We rely on our previous result [SV2] claiming that all fibres of such families are hyperkähler. This implies that the semiample Teichmüller space is covered by affine lines, which is a key step in our proof of 3.3.
@@ 2.2 Semiample line bundles and Lagrangian fibrations
Remark 2.2: The base $B$ of a Lagrangian fibration is conjectured to be biholomorphic to ${\mathbb{C}}P^{n}$. In [Mat2] Matsushita has shown that $B$ is Fano and has the same rational cohomology as ${\mathbb{C}}P^{n}$ when it is smooth. In [Hw], Hwang has shown that in this case $B$ is in fact biholomorphic to ${\mathbb{C}}P^{n}$. In [HX] Huybrechts and Xu have shown that $B$ is always biholomorphic to ${\mathbb{C}}P^{2}$ if it is normal and $\dim_{\mathbb{C}}M=4$ (see also [Ou] and [BK]).
@@
Proposition 2.3:  Let $\varphi\colon\mathcal{M}\to T$ be a smooth family of hyperkähler manifolds, $\mathcal{L}\in\operatorname{Pic}(\mathcal{M})$ a line bundle. Let $\mathcal{M}_{t}=\varphi^{-1}(t)$ and $\mathcal{L}_{t}=\mathcal{L}|_{\mathcal{M}_{t}}$ for $t\in T$. Assume that $\mathcal{L}_{t_{0}}$ is semiample for some $t_{0}\in T$ and $q(c_{1}(\mathcal{L}_{t_{0}}))=0$. Then $\mathcal{L}_{t}$ is semiample for all $t$ in some open neighbourhood of $t_{0}$.
Proof: As explained above, we may assume that there exists a Lagrangian fibration $\pi\colon\mathcal{M}_{t_{0}}\to B$ and $\mathcal{L}_{t_{0}}=\pi^{*}\mathcal{O}_{B}(1)$. We apply [Mat4, Corollary 1.3] which shows that for $t$ in a neighbourhood of $t_{0}$ the bundle $\mathcal{L}_{t}$ induces a Lagrangian fibration on $\mathcal{M}_{t}$, hence $\mathcal{L}_{t}$ is semiample.
In the sequel we will also need the following fundamental theorem about the behaviour of semiampleness in families.
Theorem 2.4:  (Lieberman–Sernesi, [LS, the main Theorem])
Let $S$ be a complex variety and $\varphi\colon\mathcal{M}\to S$ be a smooth family of complex manifolds equipped with a holomorphic line bundle $\mathcal{L}\in\operatorname{Pic}(\mathcal{M})$. Then for all $k\in{\mathbb{Z}}$ the set $\{t\in S\enskip|\enskip\kappa(\mathcal{L}_{t})\geqslant k\}$ is a countable union of complex subvarieties of $S$.
@@ 2.4 The semiample Teichmüller space
We will always study the complex structures in some fixed connected component $\mathcal{T}^{\circ}(M)$ of the Teichmüller space. It is clear that within a connected component of the semiample Teichmüller space the first Chern class of the bundle $L$ can not change. Therefore we start by fixing a primitive non-zero cohomology class $\ell\in V_{\mathbb{Z}}=H^{2}(M,\mathbb{Z})$ such that $q(\ell)=0$. If $L\in\operatorname{Pic}(M,I)$ with $c_{1}(L)=\ell$, then $\ell$ is of Hodge type $(1,1)$ on $(M,I)$, so $\mathcal{P}er(I)$ is $q$-orthogonal to $\ell$. We therefore consider the intersection
$\mathcal{D}^{\perp}_{\ell}=\mathcal{D}\cap\mathbb{P}(\ell^{\perp}).$
@@
$\mathcal{D}^{\perp}_{\ell}\to\mathcal{D}^{\prime},$
therefore the space $\mathcal{D}^{\perp}_{\ell}$ also has two connected components. The following definition makes it possible to choose one of these two connected components unambiguously.
Definition 2.6:  We define the $\ell$-period domain $\mathcal{D}_{\ell}$ to be the connected component of $\mathcal{D}^{\perp}_{\ell}$ that satisfies the following condition: for some $[I]\in\mathcal{T}^{\circ}(M)$ with $p=\mathcal{P}er(I)\in\mathcal{D}_{\ell}^{\perp}$ the class $\ell$ lies in the closure of the positive cone $V_{p}^{+}$.
@@
Proposition 2.7:  Let $M$ be a hyperkähler manifold, $\mathcal{T}^{\circ}(M,\ell)$ and $\mathcal{D}_{\ell}$ as above. Consider a a very general point $p\in\mathcal{D}_{\ell}$. Then the preimage $\mathcal{P}er^{-1}(p)$ in $\mathcal{T}^{\circ}(M,\ell)$ is a single point. Moreover, the space $\mathcal{T}^{\circ}(M,\ell)$ is connected.
@@
Proof: This statement directly follows from [Mat4] (see also 2.2). For a point $[I]\in\mathcal{T}_{\mathrm{sa}}(M,\ell)$ we consider the universal deformation $\varphi\colon\mathcal{M}\to U$ of $(M,I)$ whose base $U$ is an open neighbourhood of $[I]$ in $\mathcal{T}(M)$. The intersection $U_{\ell}=U\cap\mathcal{T}(M,\ell)$ is an open neighbourhood of $[I]$ in $\mathcal{T}(M,\ell)$, and by 2.2 after shrinking $U_{\ell}$ we get $U_{\ell}\subset\mathcal{T}_{\mathrm{sa}}(M,\ell)$.
@@
Denote by $\Gamma$ the image of $\mathcal{G}^{\circ}(M)$ in the orthogonal group $O(V_{\mathbb{Q}},q)$. Let $G$ be the stabilizer of $\ell$ in $O(V_{\mathbb{Q}},q)$. Fix a decomposition of $V_{\mathbb{Q}}$ into an orthogonal sum of two $\mathbb{Q}$-vector spaces $V_{\mathbb{Q}}=U\oplus W$, where $U$ is a 2-dimensional subspace of signature (1,1) that contains $\ell$ and $W=U^{\perp}$. The group $G$ acts on $\ell^{\perp}=W\oplus\mathbb{Q}\ell$, so it also acts on $W$. Let $K$ be the kernel of the homomorphism $G\to O(W,q)$. The group $K$ is abelian, and if we identify the Lie algebra of $G$ with $\Lambda^{2}V_{\mathbb{Q}}$, then the Lie algebra of $K$ is identified with the subspace $\ell\wedge W$. The elements of $K$ that preserve the integral structure $V_{\mathbb{Z}}$ form a lattice in $K$ that we denote by $K_{\mathbb{Z}}$. It is a free abelian group of rank $\dim W$. Let us denote by $\Gamma_{\ell}$ the intersection of $\Gamma$ with $K_{\mathbb{Z}}$. It is known that $\Gamma$ is a subgroup of finite index in $O(V_{\mathbb{Z}},q)$, see e.g. [V2]. Therefore $\Gamma_{\ell}$ is of finite index in $K_{\mathbb{Z}}$, i.e. $\Gamma_{\ell}$ is also a free abelian group of rank $\dim W$. The group $\Gamma_{\ell}$ acts on $\ell^{\perp}$ by translations in the direction of $\ell$.
Given a point $p\in\mathcal{D}_{\ell}$, the subspace $V^{\prime}_{\mathbb{C}}$ spanned by $p$, $\overline{p}$ and $\ell$ is contained in $\ell^{\perp}$. Observe that $\Gamma_{\ell}$ is contained in $K$ by construction, therefore it preserves the subspace $V^{\prime}_{\mathbb{C}}$ and its orthogonal complement. Being a subgroup of $\Gamma$, it also preserves the set of MBM classes. Hence $\Gamma_{\ell}$ preserves $\mathrm{MBM}^{\circ}_{\ell}$. The action of $\Gamma_{\ell}$ on $\ell^{\bot}$ can be given by the following formula:
$\gamma:\;x\mapsto x+q(\gamma,x)\ell,$ (2.5)
where $\gamma\in\Gamma_{\ell}={\mathbb{Z}}^{\dim W}\subset W$. Clearly, the action (2.5) has infinite orbit whenever $x\notin W^{\bot}={\mathbb{Q}}\ell$. Therefore, any element of $\mathrm{MBM}^{\circ}_{\ell}$ has an infinite $\Gamma_{\ell}$-orbit. This completes the proof.
@@ 3.1 Algebraic dimension of the manifolds with periods in $\mathcal{D}_{\ell}$
Proof: Recall that in Section 2.2 we have introduced the notation $\kappa(L)$ for the Iitaka dimension of a line bundle $L$. For a point $[I]\in\mathcal{T}^{\circ}(M,\ell)$ denote by $L_{I}$ the line bundle on $(M,I)$ with $c_{1}(L_{I})=\ell$. Define a subset $\mathcal{T}\subset\mathcal{T}^{\circ}(M,\ell)$ as follows: $[I]\in\mathcal{T}$ if and only if there exists an open neighbourhood $U\subset\mathcal{T}^{\circ}(M,\ell)$ containing $[I]$ such that for any $[I^{\prime}]\in U$ we have $\kappa(L_{I^{\prime}})\geqslant n$. By our assumptions and 2.4 $\mathcal{T}$ is a non-empty open subset of $\mathcal{T}^{\circ}(M,\ell)$. Assume that $[J]$ is a point in the closure of $\mathcal{T}$ and let $\varphi\colon\mathcal{M}\to U_{0}$ be the universal deformation of $(M,J,L_{J})$, where $U_{0}$ is a connected open neighbourhood of $[J]$. The subset $\mathcal{T}\cap U_{0}$ is non-empty and open in $U_{0}$, and for $[I]\in\mathcal{T}\cap U_{0}$ we have $\kappa(L_{I})\geqslant n$. Applying the Lieberman–Sernesi’s 2.2, we conclude that $\kappa(L_{I})\geqslant n$ for all $I\in U_{0}$. It follows that $\mathcal{T}$ is closed in $\mathcal{T}^{\circ}(M,\ell)$, and since the latter space is connected by 2.4, we have $\mathcal{T}=\mathcal{T}^{\circ}(M,\ell)$. Next we use the inequality $a(M,I)\geqslant\kappa(L_{I})$ that was recalled above to prove the first claim. The second claim now follows from [COP, Theorem 3.6].
@@
Proof: Let $p=\mathcal{P}er(I)\in\mathcal{D}_{\ell}$. By our assumption the Néron–Severi lattice of $(M,I)$ is parabolic, therefore all elements of $N^{1,1}_{p}$ are orthogonal to $\ell$. It follows that $\mathrm{MBM}^{1,1}_{p}=\mathrm{MBM}^{1,1}_{\ell,p}$ and $\mathcal{C}^{+}_{p}=\mathcal{C}^{+}_{\ell,p}$, i.e. the Kähler chambers of the positive cone coincide with the $\ell$-stable Kähler chambers. The vector $\ell$ lies in the closure of any $\ell$-stable Kähler chamber, hence $\ell$ is a nef class on $(M,I)$. The conclusion follows from 3.1 and [COP, Theorem 3.7].

### v2

@@ 1.2 The hyperkähler SYZ conjecture
In each of the currently known deformation classes of hyperkähler manifolds there exist manifolds admitting Lagrangian fibrations. Therefore a natural approach to 1.2 (at least for the known deformation classes) is to obtain the result by deformation techniques. This approach was extensively explored by Matsushita, see [Mat4], [Mat5] and [Mat6]. Using a theorem of Voisin [Voi] about deformations of Lagrangian submanifolds of a hyperkähler manifold $M$, Matsushita showed in [Mat5] that if a line bundle $L$ with $q(c_{1}(L))=0$ is semiample, then it remains semiample after a small deformation of the pair $(M,L)$, i.e. semiampleness is open in families.
In [Mat6, Theorem 1.2] Matsushita extends his results and obtains more precise statements about the behaviour of semiampleness under local deformations, assuming additionally that the base of the Lagrangian fibration defined by the semiample line bundle is isomorphic to $\mathbb{C}P^{n}$. In particular, as part of his proof [Mat6, Claim 3.2], Matsushita obtains deformation invariance of semiampleness under the assumption that the base is isomorphic to $\mathbb{C}P^{n}$.
@@ 1.3 Results
In order to prove this result we develop a moduli theory for the pairs $(M,L)$ where $M$ is a hyperkähler manifold and $L$ a semiample line bundle of vanishing BBF square on $M$. Fixing a cohomology class $\ell\in H^{2}(M,{\mathbb{Z}})$, we define a semiample Teichmüller space $\mathcal{T}_{\mathrm{sa}}(M,\ell)$ parametrizing pairs $(M,L)$ where $L$ is semiample with $c_{1}(L)=\ell$. The space $\mathcal{T}_{\mathrm{sa}}(M,\ell)$ is defined as a subset of $\mathcal{T}(M,\ell)\subset\mathcal{T}(M)$, where $\mathcal{T}(M,\ell)$ is a divisor in the usual Teichmüller space where the class $\ell$ stays of Hodge type $(1,1)$, see Section 2.4 for details. We define the period domain for $\mathcal{T}_{\mathrm{sa}}(M,\ell)$ and study the fibres of the corresponding period map. Our main result is a version of the global Torelli theorem for $\mathcal{T}_{\mathrm{sa}}(M,\ell)$, 3.3, claiming in particular that the period map is surjective when the semiample Teichmüller space is non-empty.
Our results rely only on the openness of semiampleness proved by Matsushita in [Mat5] and do not use [Mat6]. In particular, we do not assume anything about the base of our Lagrangian fibrations, which makes our results independent of the well known open problems regarding the smoothness of the base. Our main tool is the construction of special families of Lagrangian fibrations called the degenerate twistor families. We rely on our previous result [SV2] claiming that all fibres of such families are hyperkähler. This implies that the semiample Teichmüller space is covered by affine lines, which is a key step in our proof of 3.3.
@@ 2.2 Semiample line bundles and Lagrangian fibrations
Remark 2.2: The above theorem claims that all fibres of $\pi$, including the singular ones, are Lagrangian subvarieties. It means the following. Assume that $Z\subset M$ is an irreducible component of one of the fibres of $\pi$. Let $r\colon Z^{\prime}\to Z\subset M$ be a resolution of singularities. Then the theorem above claims that the dimension of $Z$ is $\frac{1}{2}\dim M$ and $r^{*}\sigma=0$, where $\sigma$ is the holomorphic symplectic form on $M$. This claim is in [Mat2, Corollary 1].
Remark 2.3: The base $B$ of a Lagrangian fibration is conjectured to be biholomorphic to ${\mathbb{C}}P^{n}$. In [Mat3] Matsushita has shown that $B$ is Fano and has the same rational cohomology as ${\mathbb{C}}P^{n}$ when it is smooth. In [Hw], Hwang has shown that in this case $B$ is in fact biholomorphic to ${\mathbb{C}}P^{n}$. In [HX] Huybrechts and Xu have shown that $B$ is always biholomorphic to ${\mathbb{C}}P^{2}$ if it is normal and $\dim_{\mathbb{C}}M=4$ (see also [Ou] and [BK]).
@@
Proposition 2.4:  (Matsushita, [Mat5, Corollary 1.3])
Let $\varphi\colon\mathcal{M}\to T$ be a smooth family of hyperkähler manifolds, $\mathcal{L}\in\operatorname{Pic}(\mathcal{M})$ a line bundle. Let $\mathcal{M}_{t}=\varphi^{-1}(t)$ and $\mathcal{L}_{t}=\mathcal{L}|_{\mathcal{M}_{t}}$ for $t\in T$. Assume that $\mathcal{L}_{t_{0}}$ is semiample for some $t_{0}\in T$ and $q(c_{1}(\mathcal{L}_{t_{0}}))=0$. Then $\mathcal{L}_{t}$ is semiample for all $t$ in some open neighbourhood of $t_{0}$.
@@ 2.4 The semiample Teichmüller space
We will always study the complex structures in some fixed connected component $\mathcal{T}^{\circ}(M)$ of the Teichmüller space. It is clear that within a connected component of the semiample Teichmüller space the first Chern class of the bundle $L$ can not change. Therefore we start by fixing a primitive non-zero cohomology class $\ell\in V_{\mathbb{Z}}=H^{2}(M,\mathbb{Z})$ such that $q(\ell)=0$. If $L\in\operatorname{Pic}(M,I)$ with $c_{1}(L)=\ell$, then $\ell$ is of Hodge type $(1,1)$ on $(M,I)$, so $\mathcal{P}er(I)$ is $q$-orthogonal to $\ell$. We need to consider only those deformations of $I$ for which $\ell$ stays of Hodge type $(1,1)$, i.e. such that $L$ deforms together with $I$ as a holomorphic line bundle. The periods of such deformations of $I$ are orthogonal to $\ell$. We therefore consider the intersection
$\widetilde{\mathcal{D}}_{\ell}=\mathcal{D}\cap\mathbb{P}(\ell^{\perp}).$
The domain $\widetilde{\mathcal{D}}_{\ell}$ has two connected components and we will consider only one of them. Below we explain this in more detail.
@@
$\widetilde{\mathcal{D}}_{\ell}\to\mathcal{D}^{\prime},$
therefore the space $\widetilde{\mathcal{D}}_{\ell}$ also has two connected components. The following definition makes it possible to choose one of these two connected components unambiguously.
Definition 2.6:  We define the $\ell$-period domain $\mathcal{D}_{\ell}$ to be the connected component of $\widetilde{\mathcal{D}}_{\ell}$ that satisfies the following condition: for some $[I]\in\mathcal{T}^{\circ}(M)$ with $p=\mathcal{P}er(I)\in\widetilde{\mathcal{D}}_{\ell}$ the class $\ell$ lies in the closure of the positive cone $V_{p}^{+}$.
@@
Proposition 2.7:  Let $M$ be a hyperkähler manifold, $\mathcal{T}^{\circ}(M,\ell)$ and $\mathcal{D}_{\ell}$ as above. Consider a very general point $p\in\mathcal{D}_{\ell}$. Then the preimage $\mathcal{P}er^{-1}(p)$ in $\mathcal{T}^{\circ}(M,\ell)$ is a single point. Moreover, the space $\mathcal{T}^{\circ}(M,\ell)$ is connected.
@@
Proof: We may assume that $\mathcal{T}_{\mathrm{sa}}(M,\ell)\neq\emptyset$, otherwise the claim clearly holds. The statement follows directly from [Mat5] (see also 2.2). For a point $[I]\in\mathcal{T}_{\mathrm{sa}}(M,\ell)$ we consider the universal deformation $\varphi\colon\mathcal{M}\to U$ of $(M,I)$ whose base $U$ is an open neighbourhood of $[I]$ in $\mathcal{T}(M)$. The intersection $U_{\ell}=U\cap\mathcal{T}(M,\ell)$ is an open neighbourhood of $[I]$ in $\mathcal{T}(M,\ell)$, and by 2.2 after shrinking $U_{\ell}$ we get $U_{\ell}\subset\mathcal{T}_{\mathrm{sa}}(M,\ell)$.
@@
Denote by $\Gamma$ the image of $\mathcal{G}^{\circ}(M)$ in the orthogonal group $O(V_{\mathbb{Q}},q)$. Let $G$ be the stabilizer of $\ell$ in $O(V_{\mathbb{Q}},q)$. Fix a decomposition of $V_{\mathbb{Q}}$ into an orthogonal sum of two $\mathbb{Q}$-vector spaces $V_{\mathbb{Q}}=U\oplus W$, where $U$ is a 2-dimensional subspace of signature (1,1) that contains $\ell$ and $W=U^{\perp}$. The group $G$ acts on $\ell^{\perp}=W\oplus\mathbb{Q}\ell$, so it also acts on $W$. Let $K$ be the kernel of the homomorphism $G\to O(W,q)$. The group $K$ is abelian, and if we identify the Lie algebra of $O(V_{\mathbb{Q}},q)$ with $\Lambda^{2}V_{\mathbb{Q}}$, then the Lie algebra of $G$ is $\Lambda^{2}(\ell^{\perp})\simeq\ell\wedge W\oplus\Lambda^{2}W$ and the Lie algebra of $K$ is $\ell\wedge W$. The elements of $K$ that preserve the integral structure $V_{\mathbb{Z}}$ form a lattice in $K$ that we denote by $K_{\mathbb{Z}}$. It is a free abelian group of rank $\dim W$. Let us denote by $\Gamma_{\ell}$ the intersection of $\Gamma$ with $K_{\mathbb{Z}}$. It is known that $\Gamma$ is a subgroup of finite index in $O(V_{\mathbb{Z}},q)$, see e.g. [V2]. Therefore $\Gamma_{\ell}$ is of finite index in $K_{\mathbb{Z}}$, i.e. $\Gamma_{\ell}$ is also a free abelian group of rank $\dim W$. The group $\Gamma_{\ell}$ acts on $\ell^{\perp}$ by translations in the direction of $\ell$. More precisely, given $y\in W$, the element $\ell\wedge y\in\ell\wedge W\simeq\mathrm{Lie}(K)$ acts on $\ell^{\perp}$ as follows:
$x\mapsto q(x,y)\ell-q(x,\ell)y=q(x,y)\ell.$
Exponentiating, we obtain the following formula for the action of $\gamma=e^{\ell\wedge y}\in K$ on $\ell^{\perp}$:
$\gamma:\;x\mapsto x+q(x,y)\ell.$ (2.5)
The subgroup $\Gamma_{\ell}$ is identified with a lattice $R\simeq{\mathbb{Z}}^{\dim W}\subset W$, so that $\gamma=e^{\ell\wedge y}\in\Gamma_{\ell}$ if and only if $y\in R$.
Given a point $p\in\mathcal{D}_{\ell}$, the subspace $V^{\prime}_{\mathbb{C}}$ spanned by $p$, $\overline{p}$ and $\ell$ is contained in $\ell^{\perp}$. Observe that $\Gamma_{\ell}$ is contained in $K$ by construction, therefore it preserves the subspace $V^{\prime}_{\mathbb{C}}$ and its orthogonal complement. Being a subgroup of $\Gamma$, it also preserves the set of MBM classes. Hence $\Gamma_{\ell}$ preserves $\mathrm{MBM}^{\circ}_{\ell}$. Clearly, the action (2.5) has an infinite orbit whenever $x\notin W^{\bot}={\mathbb{Q}}\ell$. Therefore, any element of $\mathrm{MBM}^{\circ}_{\ell}$ has an infinite $\Gamma_{\ell}$-orbit. This completes the proof.
@@ 3.1 Algebraic dimension of the manifolds with periods in $\mathcal{D}_{\ell}$
Proof: Recall that in Section 2.2 we have introduced the notation $\kappa(L)$ for the Iitaka dimension of a line bundle $L$. For a point $[I]\in\mathcal{T}^{\circ}(M,\ell)$ denote by $L_{I}$ the line bundle on $(M,I)$ with $c_{1}(L_{I})=\ell$. Define a subset $\mathcal{T}\subset\mathcal{T}^{\circ}(M,\ell)$ as follows: $[I]\in\mathcal{T}$ if and only if there exists an open neighbourhood $U\subset\mathcal{T}^{\circ}(M,\ell)$ containing $[I]$ such that for any $[I^{\prime}]\in U$ we have $\kappa(L_{I^{\prime}})\geqslant n$. By our assumptions and 2.4 $\mathcal{T}$ is a non-empty open subset of $\mathcal{T}^{\circ}(M,\ell)$. Assume that $[J]$ is a point in the closure of $\mathcal{T}$ and let $\varphi\colon\mathcal{M}\to U_{0}$ be the universal deformation of $(M,J,L_{J})$, where $U_{0}$ is a connected open neighbourhood of $[J]$. The subset $\mathcal{T}\cap U_{0}$ is non-empty and open in $U_{0}$, and for $[I]\in\mathcal{T}\cap U_{0}$ we have $\kappa(L_{I})\geqslant n$. Applying [LS, the main Theorem] to the family $\mathcal{M}$, we see that there exists an integer $k$ and a subset $W\subset U_{0}$ such that: $W$ is the complement of the union of a countable number of proper closed subvarieties of $U_{0}$, $\kappa(L_{I})\geqslant k$ for all $[I]\in U_{0}$ and $\kappa(L_{I})=k$ for $[I]\in W$. Since $\mathcal{T}\cap U_{0}$ is non-empty and open in $U_{0}$, we have $W\cap\mathcal{T}\neq\emptyset$. We conclude that $k\geqslant n$ and hence $\kappa(L_{I})\geqslant n$ for all $[I]\in U_{0}$. It follows that $\mathcal{T}$ is closed in $\mathcal{T}^{\circ}(M,\ell)$, and since the latter space is connected by 2.4, we have $\mathcal{T}=\mathcal{T}^{\circ}(M,\ell)$. Next we use the inequality $a(M,I)\geqslant\kappa(L_{I})$ that was recalled above to prove the first claim. The second claim now follows from [COP, Theorem 3.6].
@@
Proof: This is [COP, Theorem 3.4]. We give an alternative proof, since our argument will be useful below. Let $p=\mathcal{P}er(I)\in\mathcal{D}_{\ell}$. By our assumption the Néron–Severi lattice of $(M,I)$ is parabolic, therefore all elements of $N^{1,1}_{p}$ are orthogonal to $\ell$. It follows that $\mathrm{MBM}^{1,1}_{p}=\mathrm{MBM}^{1,1}_{\ell,p}$ and $\mathcal{C}^{+}_{p}=\mathcal{C}^{+}_{\ell,p}$, i.e. the Kähler chambers of the positive cone coincide with the $\ell$-stable Kähler chambers. The vector $\ell$ lies in the closure of any $\ell$-stable Kähler chamber, hence $\ell$ is a nef class on $(M,I)$. The conclusion follows from 3.1 and [COP, Theorem 3.7].